### v1

@@ 4 U2UData-2 Dataset
Figure 5: The visualization of the U2UData-2 dataset. we select two swarm UAV collaboration clips and annotate them.
Scene setting. The simulator map is first divided into 6 areas. Except for wind, which is located throughout the map, other weather is deployed in specific areas and has no intersection. For specific area locations, please watch the web page demonstration video. Since each specific area has different terrain, weather and terrain are strongly coupled. As shown in Table 3, we construct 12 scenes based on the most common weather combinations. For single-weather scenes, the trajectory of each scene collected by U2UData-2 is 5. For cross-weather scenes, the trajectory of each scene collected by U2UData-2 is 3.
Dataset collection. As shown in Table 4, the dataset is collected by 15 UAVs in autonomous formation mode for the LH task (wildlife conservation), comprising 12 scenes, 720 trajectories, and 600 seconds in length for each trajectory. The sampling interval of each sensor is 0.03 seconds and is synchronized in real time. As shown in Table 2, we collect a total of 12.96M RGB frames, 12.96M depth frames, 4.32M LiDAR frames, 25.92M airflow frames, 12.96M brightness frames, 12.96M temperature frames, 12.96M humidity frames, and 12.96M smoke frames. The total length of the entire dataset is 120 hours. The total size of U2UData-2 is 3.62T. The dataset has been open-sourced and are available for public use.
3D bounding boxes annotation. For annotating 3D bounding boxes on the gathered LiDAR data, we utilize SusTechPoint Li et al. (2020), a robust open-source labeling tool. There are a total of 15 object classes, and we annotate their 3D bounding box with 7 degrees of freedom, encompassing its location (x, y, z) and rotation (expressed as quaternions: w, x, y, z). The location (x, y, z) corresponds to the center of the bounding box. These 3D bounding boxes are annotated separately based on the global coordinate system of each UAV. This approach enables the sensor data from each UAV to be treated independently as a single-agent detection task. We initialize the relative pose of the two UAVs for each frame using positional information provided by the IMU on both UAVs.
@@
[…]
Comparison U2UData U2UData-2
Tasks Basic tasks LH tasks $\star$
@@
[…]
Table 6: Swarm UAV collaborative tracking benchmark for LH tasks in the U2UData-2 dataset.
Method AMOTA(↑) AMOTP(↑) sAMOTA(↑) MOTA(↑) MT(↑) ML(↓)
@@
U2UData-2 vs. U2UData. U2UData-2 significantly expands upon its predecessor (U2UData) by transitioning from basic collaborative perception and tracking tasks to scalable LH tasks (wildlife conservation) based on multi-UAV collaborative perception, localization, communication, navigation, tracking, and task re-allocation. As shown in Table 5, key enhancements include 40× longer UAV trajectory of each scene (600s vs 15s) and 13.7× greater total trajectory duration (120h vs 8.75h), alongside 5× increases in UAV number (15 vs 3) and tracking targets number (15 vs 3). While retaining eight sensors per UAV, U2UData-2 introduces dynamic flight algorithm selection and customizable starting points. Crucially, it adds three core innovations: a visual control window for real-time monitoring, one-click online data collection, and closed-loop algorithm validation. Those functionalities are absent in the original U2UData. Most importantly, U2UData-2 establishes a scalable framework for swarm UAV autonomous flight in dynamic open environments, which can greatly alleviate the limitations of existing datasets on algorithm development.
@@
Dataset Visualization. U2UData-2 dataset is the first large-scale swarm UAV autonomous flight dataset for the LH task (wildlife conservation). As shown in Figure 5, we select two swarm UAV collaboration clips and annotate them. The first clip ((a)-(c)) demonstrates that the target localization accuracy of a single UAV is limited due to obstacle obstruction and restricted field of view; swarm UAVs eliminate target localization errors through collaborative communication, perception, and localization. The second clip ((d)-(f)) illustrates that a single UAV makes it difficult to complete complex tasks independently due to its hardware limitations; swarm UAVs can improve the robustness of completing LH tasks through collaborative perception, communication, and task re-allocation. These visualizations highlight the dataset’s ability to design algorithms for LH tasks.
5 U2UData-2 Benchmark
@@
SOTA Algorithms. Since the algorithms of swarm UAV autonomous flight for LH tasks are still lacking, we provide a swarm UAV collaborative tracking benchmark for LH tasks in the U2UData-2 dataset. This benchmark uses 9 SOAT collaborative tracking algorithms, including No Fusion, Late Fusion, Early Fusion, When2Com Liu et al. (2020), DiscoNet Li et al. (2021), V2VNet Wang et al. (2020), V2X-ViT Xu et al. (2022b), CoBEVT Xu et al. (2022a), and Where2com Hu et al. (2022). This benchmark will be updated dynamically afterwards.
@@
Evaluation metrics. We utilize the same evaluation metrics as outlined in Weng et al. (2020) for object tracking. These metrics include: AMOTA, average multiobject tracking accuracy; AMOTP, average multiobject tracking precision; sAMOTA, scaled average multiobject tracking accuracy, which ensures a more linear representation across the entire [0, 1] range of significantly challenging tracking tasks; MOTA, multi object tracking accuracy; MT, mostly tracked trajectories; ML, mostly lost trajectories.
@@
Tracker. We’ve chosen the AB3Dmot tracker Weng et al. (2020) as our basic module of all SOAT algorithms. This tracker initially retrieves 3D object detections from a LiDAR point cloud. It subsequently integrates the 3D Kalman filter with the birth and death memory technique to guarantee efficient and resilient tracking performance. It attains state-of-the-art performance while maintaining the fastest speed.
@@
6 Conclusion
Swarm UAV autonomous flight for LH tasks is crucial. U2UData-2 is the first large-scale swarm UAV autonomous flight dataset for LH tasks and the first scalable swarm UAV data online collection and algorithm closed-loop verification platform. The dataset is captured by 15 UAVs in autonomous collaborative flight for LH tasks, comprising 12 scenes, 720 traces, 120 hours, and 600 seconds per trajectory. The platform supports the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and LH tasks. Through a visual control window, this platform allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation. U2UData-2 also provides a benchmark with 9 SOTA models. In the future, we hope U2UData-2 can assist swarm UAV algorithms in being deployed in the real world.
Appendix A U2UData-2 Platform Implementation Details
The U2UData-2 platform has been implemented with two deployment options available on Hugging Face [^11] https://huggingface.co/datasets/fengtt42/U2USim-2/tree/main.
  1. 1.
    Windows executable version: as shown in Figure 6(a), pre-compiled binary for Windows systems with integrated dependencies. There is no need to install additional packages, you can use it by double-clicking ”Landscape3.exe”. Detailed instructions are described in the README.md of the Hugging Face and the video tutorial on the paper visualization webpage [^12] https://fengtt42.github.io/U2UData-2/.
  2. 2.
    Linux executable version: as shown in Figure 6(b), optimized binary for Linux distributions supporting ROS 2 integration. There is no need to install additional packages, you can use it by running ”.Landscape3.sh” in the command line console terminal. Detailed instructions are described in the README.md of the Hugging Face.
  3. 3.
    Source code: a complete source code repository with build instructions and customization guidelines will be available on Hugging Face as soon as possible.
(a) Windows executable version
(b) Linux executable version
Figure 6: The U2UData-2 platform has been implemented with a Windows version and a Linux version.
Appendix B Visual Control Window Implementation Details
As shown in Figure 7 and Figure 8, users can collect datasets through a one-click deployment of the customized UAV starting point, weather, LH tasks, swarm UAV autonomous flight algorithms, and swarm formation mode. Users can also verify swarm UAV autonomous flight algorithms by closed-loop simulation. When running the simulation environment, you can use it by running ”python AirDroneClient.py” in the command line console terminal.
Figure 7: The source code of the visual control window.
In the visual control window [^13] https://huggingface.co/fengtt42/U2UData-2-Client/tree/main, for the platform basic capability test, users can first click the ”keyboardControl” button, and then control the UAV by clicking the following buttons: ”Take off”, ”Land”, ”Up”, ”Down”, ”Move Forward”, ”Move Backward”, ”Move Left”, and ”Move Right”. We’ve also implemented XBOX controller control. Connect your XBOX and open the simulator to control the UAV with the controller directly. It’s important to note that XBOX controller control and keyboard control are mutually exclusive.
Figure 8: U2UData-2 platform can customize the starting point-weather-task of the UAV. The wind is located throughout the map.
Appendix C Platform Visualization
As shown in Figure 9, U2UData-2 platform contains 7 weather conditions and 4 terrains. Video demos of each weather condition and images of each terrain are shown on the paper visualization webpage.
Figure 9: U2UData-2 platform contains 7 weather conditions and 4 terrains: 7 weather conditions: sunny, rain, snow, sandstorm, fog, thunder, and wind; 4 terrains: mountains, hills, plains, and basins.
Appendix D U2UData-2 Dataset
Since the entire U2UData-2 dataset is 3.6T in total, we only upload the dataset for one scenario, which is 252G in total. The data storage format is shown in Figure 10. If you need the full dataset, after the paper is accepted, you can send an email to obtain the Baidu Cloud link for downloading. PPM is a binary image format. If you need PNG format, we provide conversion code, which you can download and convert by running ”python FileConverter.py”.
Figure 10: The data storage format of U2UData-2 dataset.

### v2

@@ Introduction
Figure 4: The visualization of the U2UData+ dataset. we select two swarm UAV collaboration clips and annotate them.
Scene setting. The simulator map is first divided into 6 areas. Except for wind, which is located throughout the map, other weather is deployed in specific areas and has no intersection. For specific area locations, please watch the web page demonstration video. Since each specific area has different terrain, weather and terrain are strongly coupled. As shown in Table 3, we construct 12 scenes based on the most common weather combinations. For single-weather scenes, the trajectory of each scene collected by U2UData+ is 5. For cross-weather scenes, the trajectory of each scene collected by U2UData+ is 3.
Dataset collection. As shown in Table 4, the dataset is collected by 15 UAVs in autonomous formation mode for the ELH task (wildlife conservation), comprising 12 scenes, 720 trajectories, and 600 seconds in length for each trajectory. The sampling interval of each sensor is 0.03 seconds and is synchronized in real time. As shown in Table 2, we collect a total of 12.96M RGB frames, 12.96M depth frames, 4.32M LiDAR frames, 25.92M airflow frames, 12.96M brightness frames, 12.96M temperature frames, 12.96M humidity frames, and 12.96M smoke frames. The total length of the entire dataset is 120 hours. The total size of U2UData+ is 3.62T. The dataset has been open-sourced and are available for public use.
3D bounding boxes annotation. For annotating 3D bounding boxes on the gathered LiDAR data, we utilize SusTechPoint (SusTechPoint), a robust open-source labeling tool. There are a total of 15 object classes, and we annotate their 3D bounding box with 7 degrees of freedom, encompassing its location (x, y, z) and rotation (expressed as quaternions: w, x, y, z). The location (x, y, z) corresponds to the center of the bounding box. These 3D bounding boxes are annotated separately based on the global coordinate system of each UAV. This approach enables the sensor data from each UAV to be treated independently as a single-agent detection task. We initialize the relative pose of the two UAVs for each frame using positional information provided by the IMU on both UAVs.
@@
Table 5: A detailed comparison between U2UData and U2UData+. Basic tasks: collaborative perception and tracking. ELH tasks: wildlife conservation based on collaborative perception, localization, communication, navigation, tracking, and task re-allocation. $\star$ represents the scalable.
represents the newly added function of U2UData+.
Comparison U2UData U2UData+
Tasks Basic tasks ELH tasks $\star$
@@
Visual Control Window No
Data online Collection No
Algorithm Closed-loop No
Table 6: Swarm UAV collaborative tracking benchmark for ELH tasks in the U2UData+ dataset.
Methods AMOTA(↑) AMOTP(↑) sAMOTA(↑) MOTA(↑) MT(↑) ML(↓)
@@
U2UData+ vs. U2UData. U2UData+ significantly expands upon its predecessor (U2UData) by transitioning from basic collaborative perception and tracking tasks to scalable ELH tasks (wildlife conservation) based on multi-UAV collaborative perception, localization, communication, navigation, tracking, and task re-allocation. As shown in Table 5, key enhancements include 40× longer UAV trajectory of each scene (600s vs 15s) and 13.7× greater total trajectory duration (120h vs 8.75h), alongside 5× increases in UAV number (15 vs 3) and tracking targets number (15 vs 3). While retaining eight sensors per UAV, U2UData+ introduces dynamic flight algorithm selection and customizable starting points. Crucially, it adds three core innovations: a visual control window for real-time monitoring, one-click online data collection, and closed-loop algorithm validation. Those functionalities are absent in the original U2UData. Most importantly, U2UData+ establishes a scalable framework for swarm UAV autonomous flight in dynamic open environments, which can greatly alleviate the limitations of existing datasets on algorithm development.
@@
Dataset Visualization. U2UData+ dataset is the first large-scale swarm UAV autonomous flight dataset for the ELH task (wildlife conservation). As shown in Figure 4, we select two swarm UAV collaboration clips and annotate them. The first clip ((a)-(c)) demonstrates that the target localization accuracy of a single UAV is limited due to obstacle obstruction and restricted field of view; swarm UAVs eliminate target localization errors through collaborative communication, perception, and localization. The second clip ((d)-(f)) illustrates that a single UAV makes it difficult to complete complex tasks independently due to its hardware limitations; swarm UAVs can improve the robustness of completing ELH tasks through collaborative perception, communication, and task re-allocation. These visualizations highlight the dataset’s ability to design algorithms for ELH tasks.
U2UData+ Benchmark
@@
SOTA Algorithms. Since the algorithms of swarm UAV autonomous flight for ELH tasks are still lacking, we provide a swarm UAV collaborative tracking benchmark for ELH tasks in the U2UData+ dataset. This benchmark uses 9 SOAT collaborative tracking algorithms, including No Fusion, Late Fusion, Early Fusion, When2Com (16), DiscoNet (15), V2VNet (12), V2X-ViT (14), CoBEVT (13), and Where2com (17). This benchmark will be updated dynamically afterwards.
@@
Evaluation metrics. We utilize the same evaluation metrics as outlined in (tracking) for object tracking. These metrics include: AMOTA, average multiobject tracking accuracy; AMOTP, average multiobject tracking precision; sAMOTA, scaled average multiobject tracking accuracy, which ensures a more linear representation across the entire [0, 1] range of significantly challenging tracking tasks; MOTA, multi object tracking accuracy; MT, mostly tracked trajectories; ML, mostly lost trajectories.
@@
Tracker. We’ve chosen the AB3Dmot tracker (tracking) as our basic module of all SOAT algorithms. This tracker initially retrieves 3D object detections from a LiDAR point cloud. It subsequently integrates the 3D Kalman filter with the birth and death memory technique to guarantee efficient and resilient tracking performance. It attains state-of-the-art performance while maintaining the fastest speed.
@@
Conclusion
Swarm UAV autonomous flight for ELH tasks is crucial. U2UData+ is the first large-scale swarm UAV autonomous flight dataset for ELH tasks and the first scalable swarm UAV data online collection and algorithm closed-loop verification platform. The dataset is captured by 15 UAVs in autonomous collaborative flight for ELH tasks, comprising 12 scenes, 720 traces, 120 hours, and 600 seconds per trajectory. The platform supports the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and ELH tasks. Through a visual control window, this platform allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation. U2UData+ also provides a benchmark with 9 SOTA models. In the future, we hope U2UData+ can assist swarm UAV algorithms in being deployed in the real world.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62222209), Beijing National Research Center for Information Science and Technology under Grant No. BNR2023TD03006, China Postdoctoral Science Foundation under Grant No. 2024M751688, Postdoctoral Fellowship Program of CPSF under Grant No. GZC20240827, Open Research Fund Program of Beijing National Research Center for Information Science and Technology, and Beijing Key Lab of Networked Multimedia.